Dear all,

please find below a short summary of the activities of march 16th on the mountain related to the AO system.

March 16th

Italian people working at LBTO were:

Arcetri: A. Riccardi, R. Briguglio, M. Xompero, S. Esposito, A. Puglisi, E. Pinna, L. Fini
LBTO: G. Brusa


Total # of person: 8

AO control SW

The procedure to patch the current AOS executable at the telescope has been exercised and a few successful installations were made.

A few modifications to the AOS code were necessary to allow the patching of the current TCS build (AO2).

The configuration parameters were modified to matche the telescope environment.

A fast test showed a problem in the mirroring of variables from TCS DD to the RTDB.


LBT672a

The retroreflector holder has been removed and the mirror cover has been put on. The retroreflector for laser tracking has been installed with the help of the LBT crew.

FLAO activity

A final alignement of the AO system was attempted and was successful. The alignement was refined by using the IRTC.
The fast link fiber has been checked.
It was possible to acquire a few interaction matrices (for 60 and 200 KL modes) and
the optical loop was succwssfully closed.

Logistics

A. Riccardi, R. Briguglio, S. Esposito, A. Puglisi, E. Pinna and G. Brusa leaved Mt.Graham in the afternoon.

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback